OpenStudy (anonymous):

Cassie and Nicholas each rented a movie from two different stores. Cassie paid $4 for the video plus $1.50 for each day it was late. Nicholas spent $6 for his video plus $1.00 for each day it was late. If they both spent the same amount of money after the video was returned, how many days after the due date did they keep the video?

OpenStudy (anonymous):

if x= amount of days they were late, then: 4+1.5x=6+1*x 0.5x=6-4 x=...? how many days?

OpenStudy (anonymous):

x=2/0.5=20/5=4 x=4 days